OpenStudy (daniel.ohearn1):

Show whether the series converges or diverges using an appropriate series test or tests, Calculus B...

OpenStudy (daniel.ohearn1):

\[\sum_{n=1}^{\infty}\frac{ (-1)^nln(n)}{ n+3 }\]

OpenStudy (solomonzelman):

alternating series test

OpenStudy (solomonzelman):

You can write your series as: \(\color{#000000}{ \displaystyle \sum_{ n=1 }^{ \infty }(-1)^{n}\frac{\ln(n)}{n+3}}\)

OpenStudy (solomonzelman):

One note that (I checked) it won't be monotonically decreasing starting from n=1 (for instance |a2|<|a3|), but since \(\color{#000000}{ \displaystyle \lim_{n\to \infty}\frac{\ln(n)}{n+3}}\) therefore there is going to be some n=j starting from which \(\color{#000000}{ \displaystyle \frac{\ln(n)}{n+3}}\) is monotonically decreasing. that will show that starting from n=j, your series ___________ ? and therefore starting from n=0 the series also ________?

OpenStudy (daniel.ohearn1):

The alternating series test would be inconclusive, since Bn+1 would not be less for all n.

OpenStudy (zarkon):

for \[\sum_{n=1}^{\infty}(-1)^na_n\] all that you need is that \(a_n\) is decreasing for all n greater than some N.

OpenStudy (zarkon):

so the A.S.T. will work here

OpenStudy (daniel.ohearn1):

I'm not sure how to show that the limit would 0 for it to converge... yet.

OpenStudy (anonymous):

How about using the fact that \(\ln n<\sqrt n\) for all (positive integers) \(n\)? This comes from the fact that \(n<e^{n/2}\). Then you have \[\frac{\ln n}{n+3}<\frac{\sqrt n}{n+3}\sim\frac{1}{\sqrt n}\]

OpenStudy (daniel.ohearn1):

What is that squiggle saying?

OpenStudy (daniel.ohearn1):

It's a good idea SithsAndGiggles.

OpenStudy (daniel.ohearn1):

That comparison.

OpenStudy (daniel.ohearn1):

@SithsAndGiggles

OpenStudy (anonymous):

The \(\sim\) is meant to communicate that \(\dfrac{\sqrt n}{n+3}\) behaves like \(\dfrac{1}{\sqrt n}\), since \[\frac{\sqrt{n}}{n+3}\times\frac{\frac{1}{\sqrt n}}{\frac{1}{\sqrt n}}=\frac{1}{\sqrt n+\frac{3}{\sqrt n}}\]As \(n\to\infty\), you have that \(\dfrac{3}{\sqrt n}\to0\), and so \(\dfrac{1}{\sqrt n+\frac{3}{\sqrt n}}\) in turn behaves like \(\dfrac{1}{\sqrt n}\).

OpenStudy (zarkon):

maybe just go a step further \[\frac{\ln n}{n+3}<\frac{\sqrt n}{n+3}<\frac{\sqrt{n}}{n}=\frac{1}{\sqrt{n}}\] though this doesn't show it is a decreasing function

OpenStudy (anonymous):

Right, I assumed showing that \(b_n\to0\) was the OP's roadblock